Prev Next

SW&IT Meeting Minutes 2022-05-30

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20220530 - Bec, Matthieu 72.0
LR20220530 - Mechtley, Brandon 0.0
LR20220530 - Eychaner, Glenn 0.0
LR20220530 - Edmeades, David 0.0
LR20220530 - Grenz, Paul 0.0
LR20220530 - Hooper, Stephen 0.0
LR20220530 - Zhang, Yang 0.0
LR20220530 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports


  • Last week
    • OC, Management
    • AzCam follow up
    • SOUL reports review with BM, XZ
    • review GE's luci POI update
  • This week
    • ARGOS github synch (Lorenzo?)
    • MG support contract amends follow-up with GG
    • AO data flow, mountain synology filling up
    • OSA nightlog tool for AO? meet SA


  • Previous:
    • Usual TMS work (and discussions on core calculations, may schedule another test soon).
    • New GCS patch (John confirmed it works as intended).
    • LBC code study.
    • Met with Olga on LBC/MODS issues (we now have an LBC update list for next LBC software update, combining issues from IT8641/4566/7847 and several other things not in any IT but should be corrected/enhanced).
  • Next:
    • Vacation.


  • Last week:
    • More fixes for IT 8619 (centroid calculation); adjusted centroid algorithm, and matched RTD coordinates to Aladin coordinates.
    • Investigated IT 8640 (RTD fails to display image). Found that this problem, IT 8355 (GUI deadlocks), and IT 7493 (image subtraction problems) may all be connected to a threading problem within Aladin. Am consulting with Aladin developers and providing test cases.
  • This week:
    • Work on Observer Scheduler GUI (IT 8571, 8529).
    • Continue testing IT 8640 (RTD fails to display image) and IT 8355 (Aladin internal deadlocks).
    • Work on the three tasks at the end of the LUCI GUI document.
    • Contact J.Hill about taking on some TCS work.


  • Last week
    • SOUL-mix-mode rebased PR.
    • PR IT 8599 cube rotator
    • Updated "latest"=="stable" everywhere
      • except soul-dxwfs. Will merge Fabio's changes in when he's done .
      • dxadsec running mixed_mode_control
    • Gain ramp testing. Posted some questions to the PR.
    • Added some questions for Alfio re: getting slopes from adsec
  • This week
    • Bank A only: test plan (email aoops), test w/JCG, merge
    • Discuss with Xianyu: gain ramp questions, housekeeper
    • Add error checking code for CorrectModes?
    • Investigate "new" OffsetXY error from AOG meeting minutes
    • More GH issues.
Topic revision: r7 - 06 Jun 2022, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback